OpenStudy (meilendurcer):

Billy has 1 gallon of paint. He is going to pour it into a paint tray that measures 10 inches wide, 14 inches long, and 4 cm deep. (1 gallon = 231 in^3, 1 inch = 2.54 cm) Which of the following scenarios will happen? A. The paint will not fill the tray by 441 cm^3. B. The paint will not fill the tray by 10.53 in^3. C. The paint will fill the tray exactly. D. The paint will overfill the tray by 10.53 in^3.

OpenStudy (anonymous):

The paint tray volume is:\[10 * 14 * \left(\frac{4}{2.54}\right)\text{=}220.472 \text{ inch}^3\]The difference between the paint tray and a gallon of paint is:\[220.472-231\text{=}-10.528\]so D. is the answer, The paint will overfill the tray